# Spectral Hardening Reveals Afterglow Emergence in Long-Duration Fast X-ray Transients: A Case Study of GRB 250404A/EP250404a

### v1

@@ 4 Summary and Discussion
In this Letter, we present a detailed analysis of the long-duration emission from the fast X-ray transient EP250404a, whose gamma-ray counterpart was simultaneously detected as GRB 250404A (see Table 1). The wide-field detector EP/WXT captured the prompt emission phase of GRB 250404A, while the rapid automatic follow-up of EP/FXT recorded a continuous emission tail lasting for thousands of seconds after the trigger. We conducted a comprehensive spectral analysis of the long-duration emission in gamma-ray and X-ray bands, revealing a complete spectral evolution from the prompt emission to the afterglow. The spectral peak energy decreased from above 100 keV to below 0.5 keV within the first $\sim$ 154 s from the trigger. The spectral index $\alpha_{\rm X}$ showed a softening trend during the decaying tail before $T_{0}$ + 255 s, followed by a clear hardening, eventually reaching a plateau at $\sim$ -2. This spectral hardening coincides with a sharp flux rise at 1 keV and in the optical bands, peaking at $\sim$ $T_{0}$ + 1000 s. The subsequent decay in flux density is characterized by distinct slopes in the early and late stages. Our multiwavelength data are best fitted by a FS+RS afterglow model with extinction correction factors in the blue bands, confirming that the emerging second component at $T_{0}$ + 255 s, indicated by the spectral hardening, is the afterglow emission. These findings also provide an estimation of the prompt emission phase duration in X-rays, lasting $\sim$ 300 s, after which the X-ray emission is dominated by the afterglow.
@@
Interestingly, spectral hardening has become less uncommon in the detection of long-duration fast X-ray transients by EP, thanks to its automatic rapid follow-up capability, which enables the identification of an increasing number of events exhibiting these features. This phenomenon, characterized by the evolution of peak energy $E_{\rm p}$, softening and hardening of the spectral index $\alpha_{\rm X}$, and the appearance of spectral breaks in the SEDs, may signify a universal transition from the prompt emission phase to the afterglow phase. Benefiting from simultaneous gamma-ray observations and extensive multiwavelength follow-ups in the optical and near-infrared bands, we were able to conduct an in-depth analysis of the spectral evolution for this event. In this context, GRB 250404A/EP250404a may serve as a representative case of a specific subgroup of fast X-ray transients, distinguished by a continuous decay phase exhibiting both spectral softening and subsequent hardening. The physical mechanisms explored in this study could potentially be applicable to other similar events.
These findings suggest that the transition between the prompt emission and afterglow phases can occur in various forms, depending on the relative strengths and durations of the two components. We briefly summarize three typical scenarios (as illustrated in Figure 5) for the transition between prompt emission and afterglow in GRBs that are observed in long-duration fast X-ray transients:
@@
  • Case I: The prompt emission and afterglow components are superimposed, with the prompt emission exhibiting a harder spectral index than the afterglow at the point when the afterglow begins to dominate. In this scenario, a gradual softening of the spectral index is expected. However, the spectral overlap makes it difficult to confidently identify the afterglow component or confirm the GRB-dominated nature of the fast X-ray transient.
@@
  • Case II: The prompt and afterglow emissions are superimposed and of comparable intensity. The presence of the afterglow broadens the fast decaying profile of the prompt emission. In this case, the spectral index of the prompt emission softens beyond the value of the afterglow (we take $\sim$ $-2$ for the demonstration) before the afterglow becomes dominant. This results in a characteristic soft–hard–flat pattern in the spectral index evolution, accompanied by a smoothly decaying light curve. The spectral evolution in this case is particularly informative for identifying the emergence of the afterglow. The prompt emission and afterglow components can be disentangled by fitting the time-resolved spectra during the spectral hardening phase with two power-law models, allowing for the determination of their respective fluxes. While the total decay light curve is featureless, the decay slopes of the separated components align with standard GRB expectations: a steep decay for the prompt emission and a shallower, typical decay for the afterglow.
@@
    Figure 5: The spectral index evolutions and light curves for three scenarios of the transition from the prompt emission to the afterglow in the long-duration fast X-ray transients. The purple, blue and yellow dots represent the spectral indices of the total, prompt and afterglow emission spectra. The colored curves demonstrate the flux evolution in linear and logarithmic scale. Upper panel: Case I. Middle panel: Case II. Lower panel: Case III.
@@
  • Case III: The afterglow emerges at later times, after the prompt emission has already softened and faded. Spectral hardening is observed in the long-duration decaying tail of the continuous X-ray emission, with the flux exhibiting two distinct phases: an initial steep decay followed by a transition to a normal decay slope. This behavior is observed in this work.
@@
We are grateful to Xinglong-2.16m telescope team for the helpful discussions on the redshift determination. We also thank the support of the YNAO staff at Lijiang Observatory. We acknowledge the support by the National Key Research and Development Programs of China (2022YFF0711404, 2022SKA0130102, and 2021YFA0718500), the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos. 11833003, U2038105, U1831135, 12121003, 12393811, and 13001106), the science research grants from the China Manned Space Project with NO. CMS-CSST-2021-B11, and the Fundamental Research Funds for the Central Universities. This work is based on data obtained with the Einstein Probe, a space mission supported by the Strategic Priority Program on Space Science of the Chinese Academy of Sciences, in collaboration with ESA, MPE and CNES (Grant No. XDA15310000, No. XDA15052100). Mephisto is developed at and operated by the South-Western Institute for Astronomy Research of Yunnan University (SWIFAR-YNU), funded by the “Yunnan University Development Plan for WorldClass University” and “Yunnan University Development Plan for World-Class Astronomy Discipline”. The authors from YNU acknowledge support from the “Science & Technology Champion Project” (202005AB160002) and from two “Team Projects” - the “Innovation Team” (202105AE160021) and the “Top Team” (202305AT350002), all funded by the “Yunnan Revitalization Talent Support Program”. This work is also supported by the National Key Research and Development Program of China (2024YFA1611603) and the “Yunnan Provincial Key Laboratory of Survey Science” with project No. 202449CE340002. Y. Fang is supported by the Yunnan Fundamental Research Projects (202301AU070006). X. Wang is supported by NSFC (12288102, 12033003) and the New Cornerstone Science Foundation through the XPLORER PRIZE. Members of TNOT acknowledges financial support from the Natural Science Foundation of Xinjiang Uygur Autonomous Region under No. 2024D01D32 and the Tianshan Talent Training Program grant 2023TSYCLJ0053 and 2023TSYCCX0101, the Central Guidance for Local Science and Technology Development Fund under No. ZYYD2025QY27, the National Natural Science Foundation of China NSFC 12433007.
The SYSU 80cm infrared telescope is operated and managed by the Department of Astronomy, Sun Yat-sen University. J.M. has the financial support of the National Key R&D Program of China (2023YFE0101200), Natural Science Foundation of China 12393813, and the Yunnan Revitalization Talent Support Program (YunLing Scholar Project). AJCT acknowledges support from the Spanish Ministry project PID2023-151905OB-I00 and Junta de Andalucía grant P20_010168 and from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. DBM is funded by the European Union (ERC, HEAVYMETAL, 101071865). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140.
@@ Appendix A Spectral Fitting Results
t1 (s) t2 (s) CPL model
[…]
(keV) ($\rm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ keV^{-1}}$)

### v2

@@ 4 Summary and Discussion
In this Letter, we present a detailed analysis of the long-duration emission from the fast X-ray transient EP250404a, whose gamma-ray counterpart was simultaneously detected as GRB 250404A (see Table 1). The wide-field detector EP/WXT captured the prompt emission phase of GRB 250404A, while the rapid automatic follow-up of EP/FXT recorded a continuous emission tail lasting for thousands of seconds after the trigger. We conducted a comprehensive spectral analysis of the long-duration emission in gamma-ray and X-ray bands, revealing a complete spectral evolution from the prompt emission to the afterglow. The spectral peak energy decreased from above 100 keV to below 0.5 keV within the first $\sim$ 154 s from the trigger. The spectral index $\alpha_{\rm X}$ showed a softening trend during the decaying tail before $T_{0}$ + 255 s, followed by a clear hardening, eventually reaching a plateau at $\sim$ -2. This spectral hardening coincides with a sharp flux rise at 1 keV and in the optical bands, peaking at $\sim$ $T_{0}$ + 1000 s. The subsequent decay in flux density is characterized by distinct slopes in the early and late stages. Our multiwavelength data are best fitted by an FS+RS afterglow model with extinction correction factors in the blue bands, confirming that the emerging second component at $T_{0}$ + 255 s, indicated by the spectral hardening, is the afterglow emission. These findings also provide an estimation of the prompt emission phase duration in X-rays, lasting $\sim$ 300 s, after which the X-ray emission is dominated by the afterglow.
@@
Interestingly, spectral hardening has become less uncommon in the detection of long-duration fast X-ray transients by EP thanks to its automatic rapid follow-up capability, which enables the identification of an increasing number of events exhibiting these features. This phenomenon, characterized by the evolution of peak energy $E_{\rm p}$, softening and hardening of the spectral index $\alpha_{\rm X}$, and the appearance of spectral breaks in the SEDs, may signify a universal transition from the prompt emission phase to the afterglow phase. Benefiting from simultaneous gamma-ray observations and extensive multiwavelength follow-ups in the optical and near-infrared bands, we were able to conduct an in-depth analysis of the spectral evolution for this event. In this context, GRB 250404A/EP250404a may serve as a representative case of a specific subgroup of fast X-ray transients, distinguished by a continuous decay phase exhibiting both spectral softening and subsequent hardening. The physical mechanisms explored in this study could potentially be applicable to other similar events.
In the canonical scenario, the afterglow begins as the FS is launched into the external medium. Initially, its emission is much fainter than that of the prompt emission. As the blast wave accumulates ambient mass and the bulk Lorentz factor decreases, energy dissipation becomes more significant, and the afterglow emerges observationally, typically outshining the prompt emission at later times. Based on our findings in this case, we propose two observational indicators of the transition from prompt emission to afterglow in the X-ray band. First, a temporal break separates the steep decay (attributed to the tail of the prompt emission) from the normal decay (dominated by the afterglow). Second, the observed spectral hardening, following the initial softening associated with the prompt emission and leading into a flatter, harder afterglow spectrum, may signify a shift from prompt-dominated to afterglow-dominated emission. These findings suggest that the transition between the prompt emission and afterglow phases can occur in various forms, depending on the relative strengths and durations of the two components. We briefly summarize four typical scenarios (as illustrated in Figure 5) for the transition between prompt emission and afterglow in GRBs that are observed in long-duration fast X-ray transients.
@@
  • Case I: Smooth spectral transition with a temporal break. The prompt emission and afterglow components are superimposed, with the prompt emission exhibiting a harder spectral index than the afterglow at the point when the afterglow begins to dominate at later times. In this scenario, a gradual softening of the spectral index is expected, while the light curve exhibits a clear temporal break separating the steep decay of the prompt emission tail from the normal afterglow decay.
@@
  • Case II: Spectral hardening with smooth temporal decay. The prompt and afterglow emissions are superimposed and of comparable intensity. The presence of the afterglow broadens the fast-decaying profile of the prompt emission. In this case, the spectral index of the prompt emission softens beyond the value of the afterglow (we take $\sim$ 2-2 for the demonstration) before the afterglow becomes dominant. This results in a characteristic soft-hard-flat pattern in the spectral index evolution, accompanied by a smoothly decaying light curve. The spectral evolution in this case is particularly informative for identifying the emergence of the afterglow. The prompt emission and afterglow components can be disentangled by fitting the time-resolved spectra during the spectral hardening phase with two power-law models, allowing for the determination of their respective fluxes. While the total decay light curve is featureless, the decay slopes of the separated components align with standard GRB expectations: a steep decay for the prompt emission and a shallower, typical decay for the afterglow.
@@
    Figure 5: The spectral index evolutions and light curves for four scenarios of the transition from the prompt emission to the afterglow in the long-duration fast X-ray transients. The purple, blue, and yellow dots represent the spectral indices of the total, prompt, and afterglow emission spectra. The colored curves demonstrate the flux evolution in linear and logarithmic scale. From top to bottom: case I, case II, case III, and case IV.
@@
  • Case III: Spectral hardening with a temporal break. The afterglow dominates at later times, after the prompt emission has already softened and faded. Spectral hardening is observed in the long-duration decaying tail of the continuous X-ray emission, with the flux exhibiting two distinct phases: an initial steep decay followed by a transition to a normal decay slope. This behavior is observed in this work.
  • Case IV: Smooth spectral and temporal decay. The prompt and afterglow emissions are superimposed and have comparable intensities. At the point when the afterglow begins to dominate, the prompt emission still maintains a harder spectral index than the afterglow. Neither a significant spectral hardening nor a clear temporal break is observed. The transition appears continuous in both the light curve and spectral evolution, making it difficult to confidently identify the afterglow component or to confirm the GRB-dominated nature of the fast X-ray transient.
@@
We are grateful to Xinglong-2.16 m telescope team for the helpful discussions on the redshift determination. We are thankful for the support of the YNAO staff at Lijiang Observatory. We acknowledge the support by the National Key Research and Development Programs of China (2022YFF0711404, 2022SKA0130102, and 2021YFA0718500), the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos. 11833003, U2038105, U1831135, 12121003, 12393811, and 13001106), the science research grants from the China Manned Space Project with No. CMS-CSST-2021-B11, and the Fundamental Research Funds for the Central Universities. This work is based on data obtained with the Einstein Probe, a space mission supported by the Strategic Priority Program on Space Science of the Chinese Academy of Sciences, in collaboration with ESA, MPE, and CNES (grant No. XDA15310000, No. XDA15052100). Mephisto is developed at and operated by the South-Western Institute for Astronomy Research of Yunnan University (SWIFAR-YNU), funded by the “Yunnan University Development Plan for WorldClass University” and “Yunnan University Development Plan for World-Class Astronomy Discipline.” The authors from YNU acknowledge support from the “Science & Technology Champion Project” (202005AB160002) and from two “Team Projects,” the “Innovation Team” (202105AE160021) and the “Top Team” (202305AT350002), all funded by the “Yunnan Revitalization Talent Support Program”. This work is also supported by the National Key Research and Development Program of China (2024YFA1611603) and the “Yunnan Provincial Key Laboratory of Survey Science” with project No. 202449CE340002. Y.F. is supported by the Yunnan Fundamental Research Projects (202301AU070006). X.W. is supported by NSFC (12288102, 12033003) and the New Cornerstone Science Foundation through the XPLORER PRIZE. Members of TNOT acknowledge financial support from the Natural Science Foundation of Xinjiang Uygur Autonomous Region under No. 2024D01D32 and the Tianshan Talent Training Program grants 2023TSYCLJ0053 and 2023TSYCCX0101, the Central Guidance for Local Science and Technology Development Fund under No. ZYYD2025QY27, and the National Natural Science Foundation of China NSFC 12433007. The SYSU 80cm infrared telescope is operated and managed by the Department of Astronomy, Sun Yat-sen University. J.M. has the financial support of the National Key R&D Program of China (2023YFE0101200), Natural Science Foundation of China 12393813, and the Yunnan Revitalization Talent Support Program (YunLing Scholar Project). A.J.C.T. acknowledges support from the Spanish Ministry project PID2023-151905OB-I00 and Junta de Andalucía grant P20_010168 and from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. The 60/90 cm Schmidt telescope at the Xinglong station is operated by the WSGP group at NAOC and supported by the CAS Special Funds for Observatory and Facility Development and the Zhengjia Enterprise Group. W.L., H.Z., Y.X., W.G., L.F., and N.L. are supported by NSFC (12120101003 and 12373010), the National Key R&D Program (2022YFA1602902 and 2023YFA1607804), and the Strategic Priority Research Program of CAS (XDB0550100 and XDB0550000). D.B.M. is funded by the European Union (ERC, HEAVYMETAL, 101071865). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140.
@@ Appendix A Spectral Fitting Results
Tables 3 and 4 in Appendix A summarize the best-fitting model parameters and corresponding statistics for each time slice of the time-integrated and time-resolved spectral fittings for Fermi/GBM and EP data.
@@
$t1$ $t2$ CPL Model
$\alpha$ $E_{\rm p}$ logAA pgstat/dof BIC
(s) (s) (keV) ($\rm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ keV^{-1}}$)